OpenStudy (anonymous):

If you bought a car at $1900 which depreciates 20% each year, what will it be worth after 4 years? A.$578 B.$778 C.$988 D.$1,588

OpenStudy (anonymous):

i use y=ab^t+k formula?

OpenStudy (anonymous):

Pretty much

OpenStudy (shamil98):

Geometric Sequence! a1 = 1900 r = 0.80 n = 4

OpenStudy (anonymous):

compond intrest cormula? i use?

OpenStudy (shamil98):

\[a_n = a_1 \times r ^{n-1}\]

OpenStudy (anonymous):

ok

OpenStudy (shamil98):

\[a_4 = 1900(0.80)^{3}\]

OpenStudy (anonymous):

a 4 =972.8

OpenStudy (shamil98):

yep

OpenStudy (shamil98):

wait mb

OpenStudy (anonymous):

what do i do next?

OpenStudy (shamil98):

its B. you are looking for the 4 year so.. its n = 5 i guess. a5 = 778